OpenStudy (anonymous):

y = x + 2 2x - y = -4 Solve the system of equations using substitution. A) (-2,-4) B) (-2,0) C) (-6,-4) D) (-6,8)

OpenStudy (anonymous):

please help me:):)

OpenStudy (xapproachesinfinity):

lets label the two equation like this 1) y=x+2 2) 2x-y=-4

OpenStudy (xapproachesinfinity):

realize that the values of y is x+2 so we can replace y by x+2 in the second equation yes?

OpenStudy (xapproachesinfinity):

can you interact with me! because i don't like talking to myself

OpenStudy (xapproachesinfinity):

if you understand something say yes if not say no nope or what ever

OpenStudy (xapproachesinfinity):

so i know where do you stuck and see how to explain

OpenStudy (xapproachesinfinity):

so... do we agree that we need to replace y by x+2 in the second equation?

OpenStudy (xapproachesinfinity):

???

OpenStudy (xapproachesinfinity):

tell you are ready tag me^_-

OpenStudy (anonymous):

I think c is the answer:)

OpenStudy (xapproachesinfinity):

it is not about guessing the answer! you need to know why it is the answer

OpenStudy (xapproachesinfinity):

show how did you find it

OpenStudy (xapproachesinfinity):

I know you can do this! have trust on you abilities and you will see that you can solve this

OpenStudy (xapproachesinfinity):

come on just trust me! you can do it

OpenStudy (xapproachesinfinity):

do we have y=x+2 ? don't worry about those variables that much it just something that tells us if the right side 3 so the left hand side is also 3 it is equation because it balances the sides like scale for example y is some number and when we say y=x+2 that mean x+2 has to be the same number our goal is to find y and x that give us this equations

OpenStudy (xapproachesinfinity):

let me start over 1) y=x+2 2) 2x-y=-4 i would substitute y in the second 2) equation to get this: 2x-(x+2)=-4 (since y=x+2 this is legitimate) now i need to open the brackets 2x-x-2=-4 (x become -x and +2 becomes -2 because I'm distributing -1 if you look at this -(x+2) you see that -1 is attached to (x+2) so i need to distribute it over just like we do when we have a(b+c)=ab+ac) Now we have 2x-x-2=-4 2x-x=1x=x (if you have two things and you take out of them you are left with only one) then the equation becomes: x-2=-4 now i need to add 2 to both side x-2 =-4 +2 = +2 ---------------------- x=-4+2 now we are done just subtract 4 from 2 which mean owing since 2 is less than 4 we get x=-2 we go back to you 1) equation y=x+2 since we found x=-2 then y=-2+2 you owe me 2 dollar and i owe you 2 dollars so we are even so y=-2+2=0 y=0 summary we found x=-2 and y=0

OpenStudy (xapproachesinfinity):

so which answer do you think is right?

OpenStudy (xapproachesinfinity):

remember (x,y) is in this order x=-2 and y=0

OpenStudy (anonymous):

A

OpenStudy (xapproachesinfinity):

hmm A says x=-2 and y=-4

OpenStudy (xapproachesinfinity):

@1497

OpenStudy (anonymous):

yea I say A

OpenStudy (xapproachesinfinity):

I'm saying A in not correct since it says x=-2 and y=-4 but we found that x=-2 and y=0